# GUICourse: From General Vision Language Model to Versatile GUI Agent

### v1

@@ A.2 LLM-Auto Annotation
GUIAct (web-single): LLM-Auto Annotation.
@@
GUIAct (web-multi): Questions Acquisition.
@@
GUIChat: Question-Answer Pair Generation.
We use GPT-4 and employ a pipeline for generating both single-turn and multi-turn question-and-answer pairs. For single-turn question-and-answer generation, the process begins with generating 4 types of questions, followed by generating answers and answers with grounding. As for multi-turn question-and-answer generation, it involves generating one round of question and answer per turn based on the history of the conversation, with a maximum of five rounds. The prompts are shown in Figure 11.
Figure 11: The prompts for generating the GUIChat dataset.
@@ A.3 Crowd-sourced Annotation
Instructions Given to Participants.
We have a data annotation handbook for the annotation task for GUIAct (web-multi) data. We show the main content in Figure 12.
Cost.
(1) We pay 0.6 RMB per item to check the LLM automatic annotation samples in the GUIAct (web-single) data. The checking efficiency of each item is about 40 seconds. So, the hourly salary we pay the annotator is 54 RMB. (2) We pay 5 RMB per item to annotate GUIAct (web-multi) data, and the annotation efficiency is approximately 5 minutes per item. So, the hourly salary we pay the annotator is 60 RMB. The prices we offer are higher than the minimum hourly wage in Beijing, China (13.91 RMB).
@@ B.1 Training Details
Our GUICourse includes three datasets: GUIEnv, GUIAct, and GUIChat. (1) GUIEnv has 10M samples for pre-training (GUIEnv-global) and 0.7M samples for SFT (GUIEnv-local). Meanwhile, we split 1.8k samples in GUIEnv-local as test data for OCR and grounding abilities evaluation. (2) GUIAct is composed of three partitions: GUIAct (web-single), GUIAct (web-multi), and GUIAct (smartphone). GUIAct (web-single) is a dataset with single-step actions in website scenarios, which has about 67k instructions automatically generated by 13k screenshots. GUIAct (web-multi) is a dataset with multi-step actions and 5,696 complex instructions, which are annotated by humans. Also, we simply 9,157 instructions from AITW for smartphone scenarios, referred to as GUIAct (smartphone). There are a total of 67k, 44k, and 67k samples for the three partitions, and we split about 1.4k, 1k, and 2k samples as testing data. The residual samples are used for training. (3) GUIChat has 44k single-turn QA pairs and 6k multi-turn dialogues. We use all the samples for SFT.
@@
Our training process can be split into two stages due to the training targets. In the training Stage1, we mainly improve the OCR and text-grounding abilities of VLMs, especially in complex website scenarios. In the training Stage2, we mainly enhance the VLMs’ knowledge about GUI systems, such as the function of elements and control actions.
The whole training pipeline is fixed, but the training details are tailor-made for different VLMs, considering their uneven fundamental abilities. For Qwen-VL, which has strong fundamental abilities, we only execute the second stage of training for it. For VLMs that have weak text-grounding abilities (e.g., Fuyu-8B) and have no text-grounding ability (e.g., MiniCPM-V), we use supervised fine-tuning (using GUIEnv-local data) and pretraining (using GUIEnv-global data) methods respectively. Meanwhile, we will adjust our training stage depending on actual training objectives. For navigation-only agents, we only use the GUIAct data in the second training stage. We will mix the GUIChat data for the external functions such as the multi-turns conversation, etc.
@@
We use Type Exact Match Score (Type EM) to calculate the accuracy for predicting the names of actions, and not care about the parameters of the actions. The Type EM score is also called action type match [48] or intend match [22] in other works.
Table 7: The metrics of different actions.
@@ B.3 Case Study
Appendix C Limitations, Ethical Considerations, and Societal Impacts
Limitations.
Societal Impacts.
The screenshots in our GUIEnv dataset are collected from C4, we can’t guarantee there is no offensive content in these website screenshots.

### v2

@@ A.2.1 GUIAct (single-step website navigation)
LLM-Auto Annotation.
@@
A.2.2 GUIAct (multi-step website navigation)
Questions Acquisition.
@@ A.2.2 GUIAct (multi-step website navigation)
A.2.3 GUIChat
Question-Answer Pair Generation.
We use GPT-4 and employ a pipeline for generating both single-turn and multi-turn question-and-answer pairs. For single-turn question-and-answer generation, the process begins with generating 4 types of questions, followed by generating answers and answers with grounding. Multi-turn question-and-answer generation involves generating one round of questions and answers per turn based on the history of the conversation, with a maximum of five rounds. The prompts are shown in Figure 11.
@@ A.3 Crowd-sourced Annotation
We hire humans to help us check or annotate the GUIAct (single-step and multi-step website navigation), and GUIChat datasets.
A.3.1 GUIAct (single-step website navigation)
Data checking by human.
When we obtained auto-generated data from GPT-4V, we randomly sampled about 400 instances and discovered numerous errors, resulting in an overall accuracy of only about 55%. To address this, we contracted a professional data annotation company. We paid 0.6 RMB per instance to verify the LLM’s automatic annotations. With each instance taking approximately 40 seconds to check, the hourly wage for the annotators was 54 RMB. The entire annotation phase lasts for two months, from November 2023 to January 2024.
We have about 78k instances for review. Errors that were easy to correct, such as incorrect element IDs or action names, were adjusted accordingly. For more complex errors, such as unfiltered web pages, overall element displacement, unreasonable instructions, or mismatches between thoughts and actions, the annotators deleted the problematic samples. As a result, we ended up with 67k samples and improved the accuracy from 55% to 92%.
The checking process was divided into three batches: 7,548 samples in the first batch, 30,998 in the second, and 40,000 in the third. Each batch underwent a three-step review process: initial checking by data annotators, subsequent review by a company reviewer, and final acceptance by us. We sampled 10% of the first batch and 2% of each of the subsequent batches for quality control. Each batch was also subdivided into smaller batches (1,000 or 5,000 samples). If the accuracy in any sub-batch fell below 90%, annotators were required to recheck all data in that sub-batch until accuracy exceeded 90%. Overall, we achieved a total accuracy of approximately 92%.
A.3.2 GUIAct (multi-step website navigation)
Crowd-sourced Annotation.
We have a data annotation handbook for the annotation task for GUIAct (web-multi) data. We show the main content in Figure 12. We pay 5 RMB per item to annotate these data, and the annotation efficiency is approximately 5 minutes per item. Therefore, the hourly salary we pay the annotator is 60 RMB. The entire annotation phase lasts for four months, from September 2023 to January 2024.
A.3.3 GUIChat
Question-Answer Pair Evaluation.
In order to ensure the quality of GUIChat dataset, we evaluate by humans. The metrics we designed include "Helpfulness", "Honesty", "Harmlessness", and "Image-text correlation" four aspects: (1) Helpfulness means that the answer addresses the user’s question or request comprehensively; (2) Honesty means that the answer is factually accurate and does not have false or misleading meanings, and has roughly accurate bounding boxes; (3) Harmlessness means that the screenshots, questions, and answers do not contain offensive, discriminatory, or harmful content; (4) Image-text correlation means that the question can not be answered accurately without the screenshot. We will give a score for each aspect. If annotators think that a question-answer pair meets the requirements of a certain aspect, they will give it 1 point, otherwise 0 point. There are 44k single-turn and 6k multi-turn dialogue instructions in our GUIChat dataset. We have randomly sampled 500 question-answer pairs and evaluated them by humans. The average scores for the four aspects are $0.954$, $0.814$, $0.958$, and $0.886$ respectively. These results demonstrate that the quality of our GUIChat dataset is sufficient for training purposes. Most errors stem from screenshots of unloaded websites.
@@ B.1 Training Details
Our GUICourse includes three datasets: GUIEnv, GUIAct, and GUIChat.
(1) GUIEnv has 10M samples for pre-training (GUIEnv-global) and 0.7M samples for SFT (GUIEnv-local). Meanwhile, we split 1.8k samples in GUIEnv-local as test data for OCR and grounding abilities evaluation.
(2) GUIAct is composed of three partitions: GUIAct (web-single), GUIAct (web-multi), and GUIAct (smartphone). GUIAct (web-single) is a dataset with single-step actions in website scenarios, which has about 67k instructions automatically generated by 13k screenshots. GUIAct (web-multi) is a dataset with multi-step actions and 5,696 complex instructions, which are annotated by humans. Also, we simply 9,157 instructions from AITW for smartphone scenarios, referred to as GUIAct (smartphone). There are a total of 67k, 44k, and 67k samples for the three partitions, and we split about 1.4k, 1k, and 2k samples as testing data. The residual samples are used for SFT training.
(3) GUIChat has 44k single-turn QA pairs and 6k multi-turn dialogues. We use all the samples for SFT.
@@
Our training pipeline can be split into two stages due to the training targets. In the pre-training stage, we mainly improve the OCR and text-grounding abilities of VLMs, especially in complex website scenarios. In the SFT training stage, we mainly enhance the VLMs’ knowledge about GUI systems, such as the function of elements and control actions, as well as the conversation abilities.
We only execute the whole training pipeline based on MiniCPM-V. For other used VLMs (such as Qwen-VL and Fuyu-8B), for which we can’t acquire their pre-training data and details, we only execute the SFT training stage. (1) Pre-training stage: we mix GUIEnv-global data in the pre-training data and then execute the training. (2) SFT stage: we merge the data from three partitions as SFT data: 10% of the GUIEnv-local, all the data of GUIAct, and all the data of GUIChat. The learning rate of MiniCPM-V is 5e-7, while the learning rate of Qwen-VL and Fuyu-8B is 1e-5. All the learning scheduler types are "cosine".
@@
We use Type Exact Match Score (Type EM) to calculate the accuracy for predicting the names of actions, and not care about the parameters of the actions. The Type EM score is also called action type match (Zhan and Zhang, 2023) or intend match (Lù et al., 2024) in other works.
@@ B.3 Case Study
B.4 External Experiments
B.4.1 MiniCPM-GUI with Larger LM
We also train a larger GUI agent (MiniCPM-GUI with larger LM), similar in size to Qwen-VL, by combining GUIEnv data with its pre-training data and fine-tuning it using the GUIAct and GUIChat datasets. The new agent has better performance on our test data and outperforms SeeClick on Mind2Web, demonstrating the effectiveness of our approach when scaling up model size and incorporating comprehensive datasets. The results are shown in Figure 9 and Figure 10.
B.4.2 GUIAgent with GUIChat Data.
The main target of GUIChat data is to improve the dialogue ability of GUI agents. We have experiments that comparing the results using "GUIAct + GUIChat" datasets and "GUIAct" only based on MiniCPM-V and QWen-VL. As shown in Figure 10, the results are similar, which means the improvement of agents’ conversation ability has little impact on the action execution ability (here is the GUI navigation ability).
Agents Cross-Task Cross-Website Cross-Domain
Ele.Acc StepSR Ele.Acc StepSR Ele.Acc StepSR
SeeClick 28.3 25.5 21.4 16.4 23.2 20.8
Qwen-GUI 27.9 24.4 19.3 15.6 20.5 17.5
Fuyu-GUI 19.1 15.6 13.9 12.2 14.2 11.7
MiniCPM-GUI 23.8 20.8 20.3 17.3 17.9 14.6
+ Larger LM 36.9 33.2 34.7 29.7 35.7 31.6
Table 9: Results of our GUI agents on Mind2Web. With a larger LM, a similar size as SeeClick, MiniCPM-GUI outperforms it.
Agent Web-Single Web-Multi Smartphone Mean
GPT-4o-mini* 81.0 62.3 57.0 22.0 10.0 17.0 53.0 13.0 22.0 32.0
Fuyu-GUI 90.1 67.1 63.5 68.8 51.2 47.1 72.1 29.1 40.4 50.4
Qwen-GUI w/ GUIAct 91.9 65.6 65.0 73.1 58.0 51.2 74.7 56.8 60.8 59.0
+ GUIChat 90.9 69.4 66.7 68.9 52.5 46.8 73.0 55.7 58.1 57.2
MiniCPM-GUI w/ GUIAct 89.1 52.4 55.2 62.3 42.0 48.6 64.5 29.6 43.3 49.0
+ High Resolution 91.8 72.9 69.2 68.2 42.4 47.1 72.6 47.5 52.0 56.1
+ GUIChat 91.8 74.9 70.6 67.0 45.4 47.5 71.7 44.7 53.3 57.1
+ Larger LM 92.1 81.6 77.0 68.2 52.5 49.1 76.1 49.9 52.8 59.6
Table 10: The performance of our GUI agents on test datasets.
Figure 11: The prompts for generating the GUIChat dataset.
@@ B.4.1 MiniCPM-GUI with Larger LM
Appendix C Data Examples
We show some examples of GUICourse in this section. An example of GUIEnv-global and an example of GUIEnv-local are shown in Figure 14. Some examples of GUIAct (web-single) are shown in Figure 15, and an example of GUIAct (web-multi) is shown in Figure 16. A single-turn QA example of GUIChat is shown in Figure 17 and a multi-turn QA example of GUIChat is shown in Figure 18.
Figure 15: Four examples in the GUIAct-web-single datasets. The instructions are on the top of each screenshot, while the actions’ names are on the left-top position. Moreover, we draw a box with an index to represent the "click" action. We use green color to visualize the golden actions and use red color to visualize the predicted actions.
Figure 16: A 6-step navigation example of GUIAct-web-multi datasets. The instructions are on the top of each screenshot, while the actions’ names in each step are shown in the left-top position. These graphs show the "click", "scroll", "select_text", "copy" and "answer" actions. We use green color to visualize the golden actions in the top-6 graphs and use red color to visualize the predicted actions in the bottom-6 graphs.
Figure 17: A case of GUIChat’s single-turn QA.
Figure 18: A case of GUIChat’s multi-turn QA.
Appendix D Ethical Considerations and Societal Impacts
@@
Societal Impacts.
Our data in GUICourse don’t need personal information or offensive content. Meanwhile, we don’t select websites that might include personal information or offensive content. However, some screenshots in our GUIEnv dataset are collected from the Cleaned Common Crawl Corpus, so we cannot guarantee that these website screenshots are absolutely free of personally identifying information or offensive content.